Sixx AM Manufacturing has a target debt—equity ratio of 0.58. Its cost of equity is 16 percent, and its cost of debt is 8 percent. If the tax rate is 33 percent, what is the company's WACC?

A) 11.49%

B) 9.83%

C) 12.7%

D) 9.27%

E) 12.09%
Answer #1

Weight of equity = 1 / (1 + 0.58) = 0.6329

Weight of debt = 1 - 0.6329 = 0.3671

WACC = Weight of debt*after tax cost of debt + weight of equity*cost of equity

WACC = 0.3671*0.08*(1 - 0.33) + 0.6329*0.16

WACC = 0.01968 + 0.10126

WACC = 0.1209 or 12.09%